# Bayesian Additive Regression Trees for functional ANOVA model

### v2

@@ 7 Conclusion
ANOVA-BART can be considered as an interpretable modification of BART and shows superior performance to BART as well as other competitors on multiple benchmark datasets. Theoretically, ANOVA-BART achieves not only a (near) minimax posterior concentration rate and but also selection consistency.
There are several possible future works. First, we only consider regression problems but ANOVA-BART can be extended easily for classification problems by employing the logistic regression model. Second, theoretical results are obtained for a fixed dimension $p$ of covariates. It would be interesting to modify ANOVA-BART for high-dimensional cases where $p$ diverges also. Componentwise sparse priors would be needed for this purpose. Third, it would be useful to exploring a new MCMC algorithm for improving the scalability of ANOVA-BART.
@@
Appendix A Degree of freedom of the height vector of a binary-product tree
In this section, we explain how the height vector $\boldsymbol{\beta}=(\beta_{\mathbb{v}},\mathbb{v}\in\{-1,1\}^{|S|})$ of a binary-product tree is derived from the parameters $(S,\mathbb{s},\beta)$ and the $\mu$-identifiability condition. Proposition A.1 reveals that the degrees of freedom of $\boldsymbol{\beta}$ is 1. We also establish the boundedness of the constant associated with the heights, under the identifiability condition, as shown in Proposition A.2.
Proposition A.1.
For $S=\{\ell_{1},...,\ell_{d}\}\subseteq[p]$, $\mathbf{s}=(s_{j},j\in S)$ and $\boldsymbol{\beta}$, consider a $\mu$-identifiable binary-product tree
$\mathbb{T}^{S}(\mathbf{x}_{S})=\sum_{\mathbf{v}\in\{-1,1\}^{d}}\beta_{\mathbf{v}}\prod_{j=1}^{d}\mathbb{I}^{(v_{\ell_{j}})}(x_{\ell_{j}}-s_{\ell_{j}}>0),$
where $\mathbf{v}=(v_{j},j\in S)$ and $d\leq p$. Then, the degree of freedom of $\boldsymbol{\beta}$ is 1. Furthermore, for any $\beta\in\mathbb{R}$, we have $\beta_{\mathbf{v}}=\beta\times\mathfrak{a}_{\mathbf{v}}$, where $\mathfrak{a}_{\mathbf{v}}=\prod_{j=1}^{d}\mathfrak{a}_{\ell_{j}}^{\mathbb{I}(v_{\ell_{j}}=1)}$ and $\mathfrak{a}_{\ell_{j}}=-{\mu_{\ell_{j}}\{X_{\ell_{j}}-s_{\ell_{j}}\leq 0\}\over\mu_{\ell_{j}}\{X_{\ell_{j}}-s_{\ell_{j}}>0\}}$.
Proof.
Without loss of generality, we let $S=\{1,\ldots,d\}$. For $\mathbb{T}^{S}$, the $\mu$-identifiability condition can be rewritten as (A.1).
$\sum_{v_{j}\in\{-1,1\}^{d}}\beta_{v_{1},\ldots,v_{d}}\mu_{j}^{(v_{j})}\{X_{j}-s_{j}>0\}=0$ (A.1)
for all $(v_{k},k\neq j)\in\{-1,1\}^{d-1}$ and all $j=1,\ldots,d$.
Let $\mathcal{V}_{+}=\{j\in[d]:v_{j}=1\}$ for a given $\mathbf{v}\in\{-1,1\}^{d}$, and for given $W\subseteq\{1,\ldots,d\},$ let $\mathfrak{p}_{W}\in\{-1,1\}^{d}$ be the vector whose $i$th element is $1$ if $i\in W$ and 1-1 otherwise. For $\mathbb{v}=(-1,....,-1)$, let $\beta_{\mathbb{v}}=\beta\in\mathbb{R}$, and for $\mathbf{v}\neq(-1,...,-1)$, let $\beta_{\mathbb{v}}=\beta a_{\mathbb{v}}$, and we are going to calculate $\mathfrak{a}_{\mathbb{v}}.$
Case (1): $|\mathcal{V}_{+}|=1$. Suppose that $v_{j}=1$ for $j\in[d]$. The $\mu$-identifiability condition requires that
$\beta\mu_{j}\{X_{j}-s_{j}\leq 0\}+\beta\mathfrak{a}_{\mathfrak{p}_{j}}\mu_{j}\{X_{j}-s_{j}>0\}=0,$ (A.2)
where $\mu_{j}$ is the marginal probability measure defined in (1). Therefore, we have $\mathfrak{a}_{\mathfrak{p}_{\{j\}}}=-{\mu_{j}\{X_{j}-s_{j}\leq 0\}\over\mu_{j}\{X_{j}-s_{j}>0\}}$ and denote it as $\mathfrak{a}_{j}$.
Case (2): $|\mathcal{V}_{+}|=2$. Suppose that $v_{j}=v_{k}=1$ for $j,k\in[d]$ with $j\neq k.$ The $\mu$-identifiability condition implies
$\beta\mathfrak{a}_{j}\mu_{k}\{X_{k}-s_{k}>0\}+\beta\mathfrak{a}_{\mathfrak{p}_{\{j,k\}}}\mu_{k}\{X_{k}-s_{k}\leq 0\}=0,$
and thus we have
$\mathfrak{a}_{\mathfrak{p}_{\{j,k\}}}=-\mathfrak{a}_{j}\frac{\mu_{k}\{X_{k}-s_{k}\leq 0\}}{\mu_{k}\{X_{k}-s_{k}>0\}}=\mathfrak{a}_{j}\mathfrak{a}_{k}.$
Case (3): $|\mathcal{V}_{+}|>2$. Repeating the above computation, we conclude $\mathfrak{a}_{\mathbb{v}}=\prod_{j=1}^{d}\mathfrak{a}_{j}^{\mathbb{I}(v_{j}=1)}$ for general $\mathbb{v}.$
Proposition A.2.
If $\mu$ is the empirical distribution $\mu_{n}$, it holds that $\max_{\mathbf{v}}|\mathfrak{a}_{\mathbf{v}}|\leq n^{d}$.
Proof.
Since $1/n\leq\mu_{n,j}^{(v_{j})}\{X_{j}-s_{j}>0\}\leq 1$ for $v_{j}\in\{-1,1\}$, we have $|\mathfrak{a}_{\mathbb{v}}|\leq n^{d}$ for any $s_{j}\in\mathcal{A}_{j},$ where $\mathcal{A}_{j}$ is defined in Section 3.2. ∎
Appendix B Posterior Sampling
@@
B.1 Acceptance probability for $(S_{t},\mathbb{s}_{t})$ when $z_{t}=1$
B.1.1 Transition probability
@@
$=\frac{1}{|S_{t}^{c}|}\frac{1}{|\mathcal{A}_{j^{\text{new}}}|},$
@@
$=\frac{1}{|S_{t}|}\frac{1}{|S_{t}^{c}|}\frac{1}{|\mathcal{A}_{j^{\text{new}}}|},$
@@
$q(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}|S_{t},\mathbb{s}_{t})$ $=\frac{1}{|S_{t}^{c}|}\frac{1}{|\mathcal{A}_{j^{\text{new}}}|}\text{Pr}(\text{GROW})\mathbb{I}(|S_{t}^{\rm new}|=|S_{t}|+1)$
$+\frac{1}{|S_{t}|}\frac{1}{|S_{t}^{c}|}\frac{1}{|\mathcal{A}_{j^{\text{new}}}|}\text{Pr}(\text{CHANGE})\mathbb{I}(|S_{t}^{\text{new}}|=|S_{t}|)$
$+\frac{1}{|S_{t}|}\text{pr}(\text{PRUNE})\mathbb{I}(|S_{t}^{\text{new}}|=|S_{t}|-1).$
B.1.2 Posterior ratio
Let $r_{t,i}=y_{i}-\sum_{k\neq t}z_{k}\mathbb{T}(\mathbb{x}_{i}:S_{k},\mathbb{s}_{k},\beta_{k})$ for i=1,,ni=1,...,n. Then, conditional on $S_{t},\mathbb{s}_{t},\beta_{t}$ and $\rm others$}, $r_{t,1},\ldots,r_{t,n}$ are independent and
$r_{t,i}|S_{t},\mathbb{s}_{t},\beta_{t},{\rm others}\sim N(\mathfrak{a}_{\mathbb{v}_{t,i}}\beta_{t},\sigma^{2}),$
where $\mathbb{v}_{t,i}=(v_{t,i,j},j\in S_{t})\in\{-1,1\}^{|S_{t}|}$ such that $\prod_{j\in S_{t}}\mathbb{I}^{(v_{t,i,j})}(x_{i,j}-s_{j}>0)=1$. Since $\pi(\beta_{t})=N(0,\sigma_{\beta}^{2}),$ a straightforward calculation after integrating out $\beta_{t}$ yields that
$\pi(S_{t},\mathbb{s}_{t}|r_{1},\ldots,r_{n},{\rm others})\propto{1\over\sqrt{A}}\exp\bigg{(}{B^{2}\over 2A}\bigg{)}\pi(S_{t},\mathbb{s}_{t}),$
where
$A={1\over\sigma^{2}}\bigg{(}\sum_{i=1}^{n}\mathfrak{a}_{\mathbb{v}_{t,i}}^{2}\bigg{)}+{1\over\sigma_{\beta}^{2}}$
and
$B={1\over\sigma^{2}}\bigg{(}\sum_{i=1}^{n}\mathfrak{a}_{\mathbb{v}_{t,i}}r_{t,i}\bigg{)}.$
Thus, the ratio of the posterior of $(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new})$ and that of $(S_{t},\mathbb{s}_{t})$ is
${\pi(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}|r_{t,1},\ldots,r_{t,n},{\rm others})\over\pi(S_{t},\mathbb{s}_{t}|r_{t,1},\ldots,r_{t,n},{\rm others})}=\sqrt{{A\over A^{\rm new}}}\exp\bigg{(}{(B^{{\rm new}})^{2}\over 2A^{\rm new}}-{B^{2}\over 2A}\bigg{)}{\pi(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new})\over\pi(S_{t},\mathbb{s}_{t})}$
where $A^{\rm new}$ and $B^{\rm new}$ are defined similarly to $A$ and $B$. Moreover, the ratio of the priors is given as follows. For GROW, we have
@@ B.1.2 Posterior ratio
${\pi(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new})\over\pi(S_{t},\mathbb{s}_{t})}\times{q(S_{t},\mathbb{s}_{t}|S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new})\over q(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}|S_{t},\mathbb{s}_{t})}=1.$
Therefore, in the case of CHANGE, the prior ratio is not required.
B.1.3 Acceptance probability
@@
$p_{\text{accept}}=\min\left\{1,{{\pi(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}}|r_{t,1},\ldots,r_{t,n},{\rm others})\over\pi(S_{t},\mathbb{s}_{t}|r_{t,1},\ldots,r_{t,n},{\rm others})}{q(S_{t},\mathbb{s}_{t}|S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new})\over q(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}|S_{t},\mathbb{s}_{t})}\right\}.$
B.2 $\beta_{t}\sim\pi(\beta_{t}|S_{t},\mathbb{s}_{t},{\rm others})$ when $z_{t}=1$
Note that the distribution of $r_{t,1},\ldots,r_{t,n}$ conditional on $\beta_{t},S_{t},\mathbb{s}_{t},{\rm others}$ are independent and
$r_{t,i}|S_{t},\mathbb{s}_{t},\beta_{t},{\rm others}\sim N(\mathfrak{a}_{\mathbb{v}_{t,i}}\beta_{t},\sigma^{2}).$
Since a priori $\beta_{t}\sim N(0,\sigma_{\beta}^{2}),$ which is conjugate, it is not difficult to show that
$\beta_{t}|S_{t},\mathbb{s}_{t},{\rm others}\sim N\left({B\over A},{1\over A}\right),$
where $A={1\over\sigma^{2}}\bigg{(}\sum_{i=1}^{n}\mathfrak{a}_{\mathbb{v}_{t,i}}^{2}\bigg{)}+{1\over\sigma_{\beta}^{2}}$ and $B={1\over\sigma^{2}}\bigg{(}\sum_{i=1}^{n}\mathfrak{a}_{\mathbb{v}_{t,i}}r_{t,i}\bigg{)}.$
B.3 Sampling $\sigma^{2}|{\rm others}$
Let $r_{i}=y_{i}-\sum_{t=1}^{T_{\max}}z_{t}\mathbb{T}(\mathbb{x}_{i}:S_{t},\mathbb{s}_{t},\beta_{t})$ for i=1,,ni=1,...,n. Since $\sigma^{2}\sim IG\left({v\over 2},{v\lambda\over 2}\right)$, it follows that
$\sigma^{2}|\text{others}\sim IG\bigg{(}{v+n\over 2},{v\lambda+\sum_{i=1}^{n}r_{i}^{2}\over 2}\bigg{)}.$
B.4 Sampling $\mathbb{z}|{\rm others}$
First, we choose an index $k$ randomly from $\{1,...,T_{\max}\}$, and then let $\mathbf{z}^{\text{new}}=(z_{1}^{\text{new}},....,z_{T_{\max}}^{\text{new}})$, where $z_{k}^{\text{new}}=1-z_{k}$ and $z_{t}^{\text{new}}=z_{t}$ for $t\neq k$. Since
$y_{i}|\textbf{z}^{\text{new}},\text{others}\sim N\bigg{(}\sum_{t=1}^{T_{\max}}z_{t}^{\text{new}}\mathfrak{a}_{\mathbb{v}_{t,i}}\beta_{t},\sigma^{2}\bigg{)}$
and
$\pi\{\mathbf{z}^{\text{new}}\}$ $=\pi\bigg{\{}T=\sum_{t=1}^{T_{\max}}z_{t}^{\text{new}}\bigg{\}}\text{pr}\bigg{\{}\text{Choose}\>\>\sum_{t=1}^{T_{\max}}z_{t}^{\text{new}}\>\>\text{elements from}\>\>T_{\max}\>\>\text{elements}\bigg{\}}$
$\propto\exp\bigg{(}-C_{*}\sum_{t=1}^{T_{\max}}z_{t}^{\text{new}}\log n\bigg{)}\bigg{/}{T_{\max}\choose\sum_{t=1}^{T_{\text{max}}}z_{t}^{\text{new}}}$,
we have
$\pi\{\textbf{z}^{\text{new}}|y_{1},....,y_{n},\text{others}\}$
$\propto\exp\bigg{(}-{1\over 2\sigma^{2}}\sum_{i=1}^{n}\bigg{(}y_{i}-\sum_{t=1}^{T_{\max}}z_{t}^{\text{new}}\mathfrak{a}_{\mathbb{v}_{t,i}}\beta_{t}\bigg{)}^{2}-C_{*}\sum_{t=1}^{T_{\max}}z_{t}^{\text{new}}\log n\bigg{)}\bigg{/}{T_{\max}\choose\sum_{t=1}^{T_{\max}}z_{t}^{\text{new}}}$.
Similarly, we can calculate $\pi\{\textbf{z}|y_{1},....,y_{n},\text{others}\}$. Therefore, the posterior ratio is calculated as:
${\pi\{\textbf{z}^{\text{new}}|y_{1},....,y_{n},\text{others}\}\over\pi\{\textbf{z}|y_{1},....,y_{n},\text{others}\}}$

### v3

@@ 7 Conclusion
ANOVA-BART can be considered as an interpretable modification of BART and shows comparable performance to BART as well as other competitors on multiple benchmark data sets. Theoretically, ANOVA-BART achieves not only a (near) minimax posterior concentration rate and but also selection consistency. Moreover, the posterior concentration rate is adaptive to the interaction structure of the true function (i.e. the maximum order of signal components).
There are several possible future works. First, theoretical results are obtained for a fixed dimension $p$ of covariates. It would be interesting to modify ANOVA-BART for high-dimensional cases where $p$ diverges also. Componentwise sparse priors would be needed for this purpose. Second, it would be useful to explore a new MCMC algorithm for improving the scalability of ANOVA-BART.
@@
Appendix A Posterior Sampling
@@
A.1 Acceptance probability for $T$
For a given $\Theta=\{T,S_{1},\mathbf{s}_{1},\beta_{1},...,S_{T},\mathbf{s}_{T},\beta_{T},\psi\}$, we define a likelihood as
$\mathcal{L}(\Theta):=\prod_{i=1}^{n}p_{f(\mathbf{x}_{i}),\psi}(y_{i}),$
where $f(\mathbf{x}_{i})=\sum_{t=1}^{T}\mathbb{T}(\mathbf{x}_{i}:S_{t},\mathbf{s}_{t},\beta_{t})$.
A.1.1 $T^{\text{new}}=T+1$
Proposal for $T$.
We propose the new state
$\Theta^{\text{new}}=\{T+1,(S_{1},\mathbf{s}_{1},\beta_{1}),...,(S_{T},\mathbf{s}_{T},\beta_{T}),(S_{T+1}^{\text{new}},\mathbf{s}_{T+1}^{\text{new}},\beta_{T+1}^{\text{new}}),\psi\}$
using Random or Stepwise move.
Transition proability.
$q(\Theta|\Theta^{\text{new}})$ is given as
$q(\Theta|\Theta^{\text{new}})$
[…]
$={T^{\text{new}}\over T_{\max}}\cdot{1\over T^{\text{new}}}.$
For Random move, we have
$q(\Theta^{\text{new}}|\Theta)$ $=\text{Pr}\{\text{Increase the number of trees}\}\pi\{S_{T+1}\}\pi\{\mathbf{s}_{T+1}\}\pi\{\beta_{T+1}\}$
and for Stepwise move, we have
$q(\Theta^{\text{new}}|\Theta)$
$=\text{Pr}\{\text{Increase the number of trees}\}\text{Pr}\{\text{Choose}\>S_{T+1}\>\text{using Stepwise}\}\pi\{\mathbf{s}_{T+1}\}\pi\{\beta_{T+1}\}$
where
$\text{Pr}\{\text{Increase the number of trees}\}=1-{T^{\text{new}}/T_{\max}}$
and
$\text{Pr}\{\text{Choose}\>S_{T+1}\>\text{using Stepwise}\}$
$=\sum_{t=1}^{T}\text{Pr}\{\text{Choose}\>S_{t}\}\text{Pr}\{S_{T+1}^{\text{new}}=S_{t}\cup\{j^{\text{new}}\}\>\text{for some}\>j^{\text{new}}\in S_{t}^{c}\}$
$=\sum_{t=1}^{T}{1\over T}\mathbb{I}(\exists j^{\text{new}}\in S_{t}^{c}\>\>\text{such that}\>\>S_{t}\cup\{j^{\text{new}}\}=S_{T+1}^{\text{new}}){\mathfrak{w}_{j^{\text{new}}}\over\sum_{j\in S_{t}^{c}}\mathfrak{w}_{j}}.$
Posterior ratio.
Since the prior ratio is given as
${\pi\{\Theta^{\text{new}}\}\over\pi\{\Theta\}}={\pi\{T+1\}\pi\{S_{T+1}^{\text{new}}\}\pi\{\mathbf{s}_{T+1}^{\text{new}}\}\pi\{\beta_{T+1}^{\text{new}}\}\over\pi\{T\}},$
we have
${\pi\{\Theta^{\text{new}}|\text{others}\}\over\pi\{\Theta|\text{others}\}}={\mathcal{L}(\Theta^{\text{new}})\over\mathcal{L}(\Theta)}{\pi\{T+1\}\pi\{S_{T+1}^{\text{new}}\}\pi\{\mathbf{s}_{T+1}^{\text{new}}\}\pi\{\beta_{T+1}^{\text{new}}\}\over\pi\{T\}}$
Acceptance probability.
Note that since the dimension of the continuous parameter varies with $T$, this MH algorithm is RJMCMC. However, because the continuous parameter (height) is generated from the prior distribution, the Jacobian is equal to 1. That is, we accept the new state $\Theta^{\text{new}}$ with probability $p_{\text{accept}}$ defined as
$p_{\text{accept}}=\min\bigg\{1,{\pi\{\Theta^{\text{new}}|\text{others}\}\over\pi\{\Theta|\text{others}\}}{q(\Theta|\Theta^{\text{new}})\over q(\Theta^{\text{new}}|\Theta)}\bigg\}.$
A.1.2 $T^{\text{new}}=T-1$
As the acceptance probability for the case of $T^{\text{new}}=T-1$ can be derived by simply reversing the procedure in the case of $T^{\text{new}}=T+1$, we omit the details.
A.2 Acceptance probability for $(S_{t},\mathbb{s}_{t})$
A.2.1 Transition probability
@@
$=\frac{\mathfrak{w}_{j^{\text{new}}}}{\sum_{j\in S_{t}^{c}}\mathfrak{w}_{j}}\frac{1}{|\mathcal{A}_{j^{\text{new}}}|},$
@@
$=\text{Pr}\{\mbox{Selecting $j^{\text{new}}$ from $S_{t}^{c}$}\>\text{based on}\>P_{\boldsymbol{\mathfrak{w}}}(\cdot|S_{t}^{c}),$
@@
$=\frac{\mathfrak{w}_{j^{\text{new}}}}{\sum_{j\in S_{t}^{c}}\mathfrak{w}_{j}}\frac{1}{|S_{t}|}\frac{1}{|\mathcal{A}_{j^{\text{new}}}|},$
@@
$q(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}|S_{t},\mathbb{s}_{t})$ $=\frac{\mathfrak{w}_{j^{\text{new}}}}{\sum_{j\in S_{t}^{c}}\mathfrak{w}_{j}}\frac{1}{|\mathcal{A}_{j^{\text{new}}}|}\text{Pr}\{\text{GROW}\}\mathbb{I}(|S_{t}^{\rm new}|=|S_{t}|+1)$
$+\frac{\mathfrak{w}_{j^{\text{new}}}}{\sum_{j\in S_{t}^{c}}\mathfrak{w}_{j}}\frac{1}{|S_{t}|}\frac{1}{|\mathcal{A}_{j^{\text{new}}}|}\text{Pr}\{\text{CHANGE}\}\mathbb{I}(|S_{t}^{\text{new}}|=|S_{t}|)$
$+\frac{1}{|S_{t}|}\text{Pr}\{\text{PRUNE}\}\mathbb{I}(|S_{t}^{\text{new}}|=|S_{t}|-1).$
A.2.2 Posterior probability ratio
Let $\lambda_{t,i}=\sum_{k\neq t}\mathbb{T}(\mathbf{x}_{i}:S_{k},\mathbf{s}_{k},\beta_{k})$ for i=1,,ni=1,...,n. The likelihood ratio is given as
$\prod_{i=1}^{n}{p_{\mathbb{T}(\mathbf{x}_{i}:S_{t}^{\text{new}},\mathbf{s}_{t}^{\text{new}},\beta_{t})+\lambda_{t,i}}(y_{i})\over p_{\mathbb{T}(\mathbf{x}_{i}:S_{t},\mathbf{s}_{t},\beta_{t})+\lambda_{t,i}}(y_{i})}.$
Moreover, the ratios of the priors for GROW, PRUNE and CHANGE are given as follows. For GROW, we have
@@ A.2.2 Posterior probability ratio
${\pi\{S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}\}\over\pi\{S_{t},\mathbb{s}_{t}\}}={|\mathcal{A}_{j^{\text{deleting}}}|\over|\mathcal{A}_{j^{\text{new}}}|}.$
To sum up, the posterior probability is
${\pi\{S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}|\text{others}\}\over\pi\{S_{t},\mathbb{s}_{t}|\text{others}\}}=\bigg(\prod_{i=1}^{n}{p_{\mathbb{T}(\mathbf{x}_{i}:S_{t}^{\text{new}},\mathbf{s}_{t}^{\text{new}},\beta_{t})+\lambda_{t,i}}(y_{i})\over p_{\mathbb{T}(\mathbf{x}_{i}:S_{t},\mathbf{s}_{t},\beta_{t})+\lambda_{t,i}}(y_{i})}\bigg){\pi\{S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}\}\over\pi\{S_{t},\mathbb{s}_{t}\}}$
A.2.3 Acceptance probability
@@
$p_{\text{accept}}=\min\left\{1,{\pi\{S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}|\text{others}\}\over\pi\{S_{t},\mathbb{s}_{t}|\text{others}\}}{q(S_{t},\mathbb{s}_{t}|S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new})\over q(S_{t}^{\rm new},\mathbb{s}_{t}^{\rm new}|S_{t},\mathbb{s}_{t})}\right\}.$
A.3 Acceptance probability for $\beta_{t}$ and $\mathbf{s}_{t}$
A.3.1 Proposal for $\beta_{t}$
To propose new state $\beta_{t}^{\text{new}}$, we use Langevin Dynamics, i.e.,
$\beta_{t}^{\text{new}}=\beta_{t}+{\varepsilon^{2}\over 2}{\partial\over\partial\beta_{t}}\log\pi\{\beta_{t}|\text{others}\}+\varepsilon m,$
where $m\sim N(0,1)$, $\varepsilon$ is a step size and ${\partial\over\partial\beta_{t}}\log\pi\{\beta_{t}|\text{others}\}$ is defined as